OpenStudy (anonymous):

y varies inversely as twice x. When x = 9, y = 2. Find y when x = 3.

OpenStudy (anonymous):

y varies inversely as twice x means y = k * (1/2x) Put x = 9, y = 2 in y = k *( 1/2x) 2=k * (1/(2*9)) so k=18*2=36 To find y when x = 3 so putting x=3 and k= 36 y=36/(2*3) =36/6 =6

OpenStudy (anonymous):

thanks u